OpenStudy (anonymous):

What is the exponential function modeled by the following table? x f(x) 3 8 4 16 5 32 f(x) = 2x f(x) = x2 f(x) = x3 f(x) = 3x

OpenStudy (anonymous):

to me it looks more like \[f(x)=2^x\]

OpenStudy (anonymous):

Could you explain? please

OpenStudy (anonymous):

2^3=8 2^4=16 and 2^5=32

OpenStudy (anonymous):

ok great! thanks

OpenStudy (anonymous):

your welcome :)